OpenStudy (anonymous):

Divide. –63 ÷ (–9 ) = A. –9 B. –7 C. 7 D. 9

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Remember when dividing or multiplying a negative by a negative you will get a positive.

OpenStudy (anonymous):

maybe C also can u help with 4 more

OpenStudy (anonymous):

@Loopki40

OpenStudy (anonymous):

Yes C is correct! and sure I can help with more.

OpenStudy (anonymous):

Divide. – 5.4 ÷ 0.9 = A. –6.0 B. –0.6 C. 0.6 D. 6.0

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Remember you are dividing a negative by a positive. So your answer will be positive.

OpenStudy (anonymous):

won't**

OpenStudy (anonymous):

so A? i know its 1 of them

OpenStudy (anonymous):

yes A

OpenStudy (anonymous):

Divide. – 5.4 ÷ 0.9 = A. –6.0 B. –0.6 C. 0.6 D. 6.0

OpenStudy (anonymous):

not that 1

OpenStudy (anonymous):

Simplify. 24 ÷ (–8 – (–4)) A. –6 B. –2 C. 2 D. 6

OpenStudy (anonymous):

Work from the inside out. So do the (-8 - (-4)) first. Remember, when you have two minus signs that really means to add. So (-8 - (-4)) can be rewritten as (-8 + 4)

OpenStudy (anonymous):

Then -8 + 4 can be rewritten as 4 - 8. So your problem transforms into 24/-4

OpenStudy (anonymous):

B or C for sure

OpenStudy (anonymous):

towards B

OpenStudy (anonymous):

24/-4 is essentially splitting 24 into 4 parts. What number multiplied by 4 gives 24? But! the 24 is being divided by a negative number, which means the answer is negative. So to be more specific what Number multiplied by -4 gives 24.

OpenStudy (anonymous):

it would be C because when multiplying or dividing if sighs are the same its positive different signs negative

OpenStudy (anonymous):

can u help with 2 more

OpenStudy (anonymous):

The answer is not B or C.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

explain how it is C completely

OpenStudy (anonymous):

It is not C. Or B. You are dividing 24 by -4. Shotgun please think of what number goes into 24 exactly 4 times. The answer is -6. Because -6*4 = 24.

OpenStudy (anonymous):

-24*

OpenStudy (anonymous):

I AM TALKING ABOUT -63/(-9)

OpenStudy (anonymous):

Oh... xD that problem was already solved. :P

OpenStudy (anonymous):

OH OKAY LOL

OpenStudy (anonymous):

Simplify. 24 ÷ (–8 – (–4)) A. –6 B. –2 C. 2 D. 6

OpenStudy (anonymous):

@pooja195